OpenStudy (anonymous):

Solve and describe the steps you took to solve: 12+3(m+1)=-2+4(m+6)

OpenStudy (mathstudent55):

@22Megan22 Do you know the distributive property?

OpenStudy (anonymous):

use the distributive property then its a two step equation

OpenStudy (rina.r):

12+3m+3=-2+4m+24 collect the like terms 12+2+3-24=+4m-3m 17-24 = 1m -7 =1m m=7/1 =7

OpenStudy (mathstudent55):

@Rina.r You're almost correct. This line is correct: -7 =1m m=-7/1 ^ you forgot the 7 was negative =-7 The answer is m = -7.

OpenStudy (anonymous):

Hmm, ok

OpenStudy (rina.r):

yeah, same mistake i make in exams:(

OpenStudy (mathstudent55):

Ther's a way of catching these mistakes. Take the solution and plug it into the original equation. If it works it's correct. For m = 7, you get: 12+3(7+1)=-2+4(7+6) 12 + 3*8 = -2 + 4*13 12 + 24 = -2 + 52 36 = 50 which is false, so m = 7 is not correct. For m = -7, we get: 12+3(-7+1)=-2+4(-7+6) 12 + 3*(-6) = -2 + 4*(-1) 12 - 18 = -2 + (-4) -6 = -6 which is true, so m = -7 is correct.

OpenStudy (rina.r):

Thanks